Circular No. 6804

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE P. Nugent, Lawrence Berkeley National Laboratory (LBNL), on behalf

of the Supernova Cosmology Project [cf. IAUC 6596, plus G. Aldering, P. Castro, N. Nunes, and R. Quimby (LBNL)], reports the discovery of seventeen supernovae, all discovered near maximum light and found with the Cerro Tololo 4-m telescope in their High Redshift Supernova Search:

SN  1997 UT  α_{2000}  δ_{2000}  R  z type
1997ek Dec. 28 4^{h}56^{m}11.^{s}63 − 3^{o}41′26′′.0 23.8 0.86 Ia
1997el Dec. 28 4 56 41.21 − 3 27 54.7 23.1 0.64 Ia
1997em Dec. 28 4 56 50.44 − 3 51 37.1 23.6 0.46 Ia
1997en Dec. 28 4 56 57.21 − 4 11 46.0 24.4 0.77 Ia?
1997eo Dec. 28 4 57 32.35 − 3 30 06.2 24.5 0.70 Ia?
1997ep Dec. 28 4 57 48.59 − 3 42 44.7 22.4 0.46 Ia
1997eq Dec. 28 4 58 56.32 − 3 59 29.4 22.5 0.54 Ia
1997er Dec. 28 5 00 38.56 − 3 59 32.2 22.3 0.47 Ia
1997es Dec. 28 8 18 40.65 + 3 13 36.5 24.3 0.65 Ia?
1997et Dec. 28 8 22 53.73 + 3 52 14.7 23.2 0.63 Ia
1997eu Dec. 28 8 23 00.45 + 4 08 27.3 22.4 0.59 Ia
1997ev Dec. 28 8 24 20.28 + 3 51 36.0 23.0 0.43 II?
1997ew Dec. 28 8 24 25.09 + 3 49 08.0 23.9 0.59 II/Ic?
1997ex Dec. 28 8 24 27.87 + 3 52 05.8 21.6 0.36 Ia
1997ey Dec. 29 4 56 58.19 − 2 37 36.7 22.9 0.58 Ia
1997ez Dec. 29 8 21 38.13 + 3 25 10.5 23.4 0.78 Ia
1997fa Dec. 29 8 22 03.72 + 3 24 24.9 22.5 0.50 Ia

The supernova spectra were obtained at Mauna Kea (Keck telescope) on 1997 Dec. 31, 1998 Jan. 1, and 2. The redshifts were determined from hostgalaxy emission and absorption lines, and all the supernovae were found to be within 2′′.5 of the host-galaxy center. R- and I-band photometry of the lightcurves are being obtained for all supernovae from groundbased telescopes, while R-, I- and J-band photometry for SNe 1997ek, 1997es, 1997ep, and 1997ez are being obtained with the Hubble Space Telescope; finding charts are available from saul@lbl.gov.

COMET 104P/KOWAL 2 Visual m_{1} estimates by M. Lehký, Hradec Králové, Czech Republic

(0.42-m reflector): 1997 Oct. 19.78 UT, 14.3; Nov. 1.83, 13.9; 20.75, 14.1; Dec. 16.74, 13.8; 21.74, 13.9; 30.75, 13.6; 31.73, 13.7; 1998 Jan. 1.74, 13.5.

1998 January 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green